# seq-JEPA: Autoregressive Predictive Learning of Invariant-Equivariant World Models

### v1

@@ 4.1 Compared methods and baselines
We compare seq-JEPA’s performance with both invariant and equivariant SSL methods. For invariant methods, we compare with SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and VICReg (Bardes et al., 2022). For equivariant methods, we compare with SEN (Park et al., 2022), EquiMod (Devillers & Lefort, 2022), SIE (Garrido et al., 2023), and ContextSSL (Gupta et al., 2024). Additionally, we consider another equivariant baseline which we call Conditional BYOL, where BYOL’s predictor is conditioned on the action that transforms the view of the target branch to the view of the online branch, with the aim that transformation information is also encoded in representations. This architecture is equivalent to a two-view version of seq-JEPA with a sequence length of one during both training and inference, and with no sequence aggregator on top of representations. Furthermore, as a baseline for predictive learning across saccades, we consider a model which we call Conv-JEPA. This baseline has the same architecture as Conditional BYOL, and to train it, we consider the same sequence of patches that seq-JEPA uses (sampled from image saliencies), predict the representation of the last patch from each individual patch in the sequence, and sum their prediction losses. For architectural details of each baseline, see Section A.3.
@@ 4.2 Training protocol
For all methods, we use a ResNet-18 (He et al., 2016) (not pretrained) as the backbone encoder. For experiments that use CIFAR100 and low-resolution STL-10 patches in predictive learning across saccades, we use the CIFAR variant of ResNet-18. For action conditioning, similar to EquiMod (Devillers & Lefort, 2022), we use a learnable linear projection to learn action embeddings. In our default setting for all experiments with action conditioning, the action projector is 128-dimensional. For the sequence aggregator, we use a small transformer encoder (Vaswani et al., 2017) with three layers, four attention heads, and post-normalization. For the predictor, we use an MLP with a hidden layer of 1024 dimensions and ReLU activation. In baseline comparisons, in order to control for and eliminate any performance gain resulting from using a transformer encoder in seq-JEPA instead of an MLP projection head, we train two variants of baselines that typically employ an MLP projector: one maintaining the original MLP projector and another where the projector is replaced with our transformer encoder. We report the higher performance among the two models. For a fair comparison, all models are trained for a fixed batch size of 512 and the same number of epochs. For 3DIEBench experiments, all models are trained for 1000 epochs, while for other datasets, they are trained for 2000 epochs. For models that use a transformer as projector, including seq-JEPA, we use the AdamW optimizer with default $\beta_{1}$ and $\beta_{2}$, a weight decay of $0.001$, and a learning rate of $4\times 10^{-4}$ with a linear warmup for $20$ epochs starting from $10^{-5}$ followed by a cosine decay back to $10^{-5}$. For models that use an MLP projector, we use the Adam optimizer with a learning rate of $10^{-3}$, default $\beta_{1}$ and $\beta_{2}$, and no weight decay.
@@ 4.3 Evaluation metrics and protocol
To measure equivariance with respect to a specific transformation, we perform action decoding from encoder representations as in Garrido et al. (2023). Specifically, given two transformed views, we extract representations from the frozen encoder, concatenate them, and train a regression head on top to predict the relative transformation (action) between the two. We report the $R^{2}$ score of this regressor on the test set as the equivariance metric. Although not always a perfect measure of invariance (Wang et al., 2024), we report top-1 classification accuracy of a linear probe as an approximate metric for representational invariance. For all baselines, we report the accuracy based on encoder representations. For seq-JEPA, we measure accuracies on top of the aggregate representation ($z_{AGG}$ in Figure 2). When reporting accuracy based on $z_{AGG}$, we report the number of observation views (context lengths) used during both training and inference. For completeness, we also report classification performance on top of encoder representation ($z_{i}$ in Figure 2) for all seq-JEPA models in Section B.4. Details on training the evaluation probes are given in Section A.2.
@@ 5.1 Quantitative evaluation on 3DIEBench
We use the 3DIEBench benchmark to quantitatively compare invariant and equivariant representations in seq-JEPA with baseline methods. This benchmark allows us to measure equivariance through decoding 3D object rotations while enabling invariance measurement through object category classification. Table 1 provides a summary of our evaluation results on the 3DIEBench dataset where equivariant methods have been conditioned on rotation. In addition to the relative rotation between two views, in the last column, we provide the $R^{2}$ score for predicting individual transformation parameters from representations of a single view. For seq-JEPA, we trained models with training sequence lengths between one to four (denoted by $M_{tr}$ in the table) and measure the linear classification performance with different inference lengths (denoted by $M_{val}$) on top of the aggregate representation. Among the invariant methods, BYOL achieves the highest classification accuracy, yet does not offer a high level of equivariance. Among equivariant baselines, while EquiMod and SEN offer even better classification accuracy compared to invariant methods, in terms of equivariance, they lag behind SIE and ContextSSL which provide the best rotation prediction performance due to their specialized equivariance predictors and loss functions. Even though both SIE and ContextSSL representations become highly sensitive to rotation, they do not fare well in linear probe classification. Seq-JEPA is able to achieve a rotation $R^{2}$ on a par with the best performing equivariant method while outperforming all methods in classification thanks to sequence aggregation during training and inference with a margin that increases with the inference sequence length. Finally, when action conditioning is ablated (last row in the table), seq-JEPA loses its equivariant properties, yet maintains a strong classification performance. For additional results on 3DIEBench including models with different training and inference sequence lengths and models conditioned on rotation and color, see Sections B.1 to B.3.
Table 1: Quantitative evaluation of learned representations on 3DIEBench for rotation prediction (equivariance) and linear probe classification (invariance). Equivariant models and seq-JEPA are conditioned on rotation. For seq-JEPA, the training sequence length is denoted by $M_{tr}$ and probing on top of $z_{AGG}$ is done with an inference length $M_{val}$.
Method Classification (top-1) Relative Rotation ($R^{2}$) Individual Rotation ($R^{2}$)
@@
seq-JEPA ($M_{tr}=1$, $M_{val}=1$) 84.08 0.65 0.69
seq-JEPA ($M_{tr}=1$, $M_{val}=3$) 85.31 0.65 0.69
seq-JEPA ($M_{tr}=3$, $M_{val}=3$) 86.14 0.71 0.74
seq-JEPA ($M_{tr}=3$, $M_{val}=5$) 87.41 0.71 0.74
seq-JEPA (w/o act cond, $M_{tr,val}=3$) 86.05 0.29 0.37
@@ 5.2 Qualitative evaluation on 3DIEBench
In order to qualitatively evaluate the representations learned by seq-JEPA and shed some light on its invariance-equivariance properties, we conduct two different visualizations. First, to see equivariance in representational space, we retrieve the three nearest representations of a query image from the validation set of 3DIEBench (Figure 4). All methods successfully retrieve the correct object category, while only seq-JEPA and SIE, which demonstrate high quantitative equivariance performance (Table 1), maintain rotation equivariance across all three nearest representations.
Figure 4: Retrieval of nearest representations; given a query image, we extract the three nearest encoder representations in the validation set of 3DIEBench. The retrieved views of models with the highest quantitative rotation equivariance performance maintain the rotation of the query image.
@@
In the second visualization, we plot the 2D UMAP projections of seq-JEPA’s encoder and aggregate representations (Figure 5) for the model trained conditioned on rotation. The left panel of of Figure 5 shows encoder representations colored by class label, while the middle panel displays the same encoder representations colored by rotation angle. The smooth color gradation across the map within each class cluster in this middle panel suggests that the encoder captures rotation angle as a continuous factor, implying equivariance to rotation (e.g., the red class in the bottom-right corner of the right panel and the corresponding part in the middle panel). The right panel shows aggregate representations colored by class label. Comparing the class-colored plots (left and right panels), we observe that both encoder and aggregate representations contain class information. However, when we aggregate multiple views of a sample, some of the intra-class variability (resulting from transformations such as rotation) is eliminated, causing each class’s representational cluster to become more homogeneous. This aggregation procedure likely reduces variation due to rotation and makes the representations more invariant, resulting in decreased intra-class spread and increased inter-class distance. We create a similar UMAP visualization for seq-JEPA with ablated rotation conditioning in Section B.5 which highlights the role of action conditioning in achieving equivariance to rotation.
@@
Figure 5: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations. The model is trained on 3DIEBench and conditioned on rotation with $M_{tr}=3$ and $M_{val}=5$. Encoder representations for each view observation, color-coded by class category (left) and rotation angle (middle). The aggregate token representation for $M_{val}=5$ observations during inference, color-coded by class category (right).
5.3 Evaluation of models trained using hand-crafted augmentations
@@
We next quantitatively assess invariance and equivariance for the second set of transformations, i.e. hand-crafted augmentations. Table 2 provides a summary of our evaluation on models trained on CIFAR100 and Tiny ImageNet. For each equivariant method and seq-JEPA, we trained four different variations. In three variations, the model is conditioned on only one of the augmentations among cropping, color jittering, and blurring, and in the last variation, on all three augmentations. Considering equivariance to augmentations, seq-JEPA outperforms invariant and equivariant baselines by a large margin. Notably, except for the model trained on CIFAR-100 and conditioned on blur, the best equivariance performance for a given augmentation is achieved when the model is specialized and conditioned only on that augmentation. Furthermore, ablating actions (last row in the table) causes seq-JEPA to achieve a lower equivariance performance across transformations compared to action-conditioned models. In terms of linear probe classification accuracy on top of aggregate representations, our framework is competitive with both invariant and equivariant families, yet does not outperform all of them. For additional results with different training and inference sequence lengths for seq-JEPA models trained with augmented views, see Section B.3.
@@
Table 2: Quantitative evaluation of learned representations with hand-crafted augmentations on CIFAR100 and Tiny ImageNet; equivariance is measured by predicting relative transformation parameters associated with crop, color jitter, or blur augmentations. For seq-JEPA, the training sequence length is denoted by $M_{tr}$ and probing is done with an inference length ($M_{val}$) of five.
@@
5.4 Predictive learning across saccades
Quantitative evaluation. In our third action-observation setting, we consider predictive learning across simulated eye movements to exhibit seq-JEPA’s ability in leveraging a sequence of partial observations alongside actions to learn visual representations. Table 3 summarizes our evaluation results for predictive learning across saccades on STL-10. While seq-JEPA only uses low-resolution patches with no augmentation, it achieves a classification performance comparable to SimCLR, an invariant model that is trained using strongly augmented image views with full resolution. Ablating action conditioning causes the accuracy on top of the aggregate representations to drop significantly, indicating that positional awareness is essential to forming semantic representations across simulated eye movements. Furthermore, comparison with the Conv-JEPA baseline highlights the importance of aggregating a sequence of patch representations in contrast to simply accumulating prediction errors between pairs of patches. Additionally, the last two ablation results in the table show that saliency-driven sampling of patches alongside IoR play a key role in sampling informative and non-overlapping patches and achieving high-quality aggregate representations. Interestingly, while random uniform patch sampling negatively impacts classification accuracy, it results in the highest positional equivariance, as the model samples patches and corresponding saccade actions from across the entire image and not just the salient regions.
Table 3: Quantitative evaluation of predictive learning across saccades on STL-10. Equivariance is measured with respect to the position of a fixated patch on the image. Seq-JEPA models use only low-resolution patches with no augmentation during training and inference. Unless stated otherwise, training and inference lengths ($M_{tr}$ and $M_{val}$) are set to four for all seq-JEPA variants.
Conditioning Classification (top-1) Position ($R^{2}$)
@@ 5.4 Predictive learning across saccades
position seq-JEPA (no saliency) 79.85 0.88
position seq-JEPA (no IoR) 77.97 0.85
Qualitative evaluation. Figure 6 shows UMAP projections for seq-JEPA representations trained on STL-10 via predictive learning across saccades. As for rotation in Figure 5, equivariance to fixation position can be observed from the bottom panels that show UMAP projection of encoder representations color-coded by X and Y coordinates. Similar to rotation in 3DIEBench, we provide a similar UMAP visualization for seq-JEPA with ablated position conditioning in Section B.5 for predictive learning across saccades to highlight the role of action conditioning.
Figure 6: Visualization of 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations. The model is trained on STL-10 via predictive learning across saccades and conditioned on fixation position with $M_{tr,val}=4$. Encoder representations for each view observation, color-coded by class category (top-left) and $X$ and $Y$ pixel coordinates (bottom panels). The aggregate token representation for $M_{val}=4$ observation patches during inference, color-coded by class category (top-right).
5.5 Path integration across actions
While an agent executes a sequence of actions in an environment, transitioning from an initial state to a final state, it should be capable of tracking its position by integrating its own actions. This is also a crucial cognitive ability that enables animals to estimate their current state in their habitat (McNaughton et al., 2006). Here, we evaluate whether pre-trained seq-JEPA is able to solve path integration. Given the sequence of observations $\{x_{i}\}_{i=1}^{M+1}$ generated from transformations $\{t_{i}\}_{i=1}^{M+1}$ and the corresponding action embeddings $\{a_{i}\}_{i=1}^{M}$, we define the task of path integration over the sequence of actions as predicting the relative action that would directly transform $x_{1}$ to $x_{M+1}$ (which we denote as $a_{1,M}$) given $z_{AGG}$ and $a_{M}$. In other words, given the aggregate representation of a sequence of action-observation pairs and the next action, we would like to predict the overall position change from the starting point ($x_{1}$) to the end point ($x_{M+1}$). We consider path integration for rotation angles in 3DIEBench and across eye movements with STL-10. For rotations, the task is integrating a series of object rotations from the first view to the last, i.e. angular path integration. For eye movements, the task is integrating the eye movements from the first saccade to the last, i.e. visual path integration. To measure path integration performance, we train a regression head on top of the concatenation of $z_{AGG}$ and $a_{M}$ to predict $a_{1,M}$ for a given inference sequence length. Figure 7 shows the results for both angular and visual path integration. The red curve corresponds to the performance of the original seq-JEPA. The blue curve corresponds to experiments in which the action embeddings are ablated (they are set to zero during inference for all views). The green curve corresponds to experiments in which the encoder (visual) representations are ablated (they are set to zero when being concatenated with actions before being fed to the transformer). As expected, path integration becomes more difficult as the number of observations increases (red curves). Ablating action conditioning (blue curves) results in failure of path integration. On the other hand, ablating the visual representations (green curves) results only in a small performance drop compared to the original model, indicating that action conditioning is the key factor enabling path integration.
@@ 5.5 Path integration across actions
5.6 Role of action conditioning
To better understand the importance of action conditioning in learning highly equivariant representations, we perform ablation experiments for seq-JEPA trained on 3DIEBench by changing the dimensionality of action embeddings and ablating actions from one of the two or both components of seq-JEPA, i.e. the transformer encoder and the MLP predictor. These ablation results are provided in Table 4. An important observation is that both predictor and transformer action conditioning are important in achieving the best equivariant performance, yet predictor action conditioning plays a more important role. Furthermore, learnable action embeddings play a key role in developing equivariance to rotation, although this equivariance shows little sensitivity to the dimensionality of the action embedding. Additionally, classification performance is slightly boosted in the presence of action conditioning, yet its role in this case is not as important compared to positional embeddings in predictive learning across saccades.
Table 4: Ablation results on the effect of seq-JEPA’s action conditioning trained on 3DIEBench (equivariance to rotation); we investigate the effects of partial or complete action ablations and changing the dimensionality of action embeddings on linear classification and rotation prediction performance. All models have $M_{tr}=3$ and $M_{val}=5$ and results are reported over three randoms seeds.
Variant Classification (top-1) Rotation ($R^{2}$)

### v2

@@ 3.1 Compared methods and baselines
We compare seq-JEPA against both invariant and equivariant SSL baselines. Invariant methods include SimCLR (Chen et al., 2020), BYOL (Grill et al., 2020), and VICReg (Bardes et al., 2022). Equivariant methods include SEN (Park et al., 2022), EquiMod (Devillers and Lefort, 2022), SIE (Garrido et al., 2023), and ContextSSL (Gupta et al., 2024). For all baselines, architectural details are given in Appendix A.3. We also evaluate two hybrid baselines based on our architecture:
  • Conditional BYOL. A two-view version of seq-JEPA with no sequence aggregator, where BYOL’s predictor is conditioned on the relative transformation between target and online views. This encourages representations to encode transformation information.
  • Conv-JEPA. A baseline for the saccades setting. It uses the same sequence of saliency-sampled patches as seq-JEPA and predicts the final patch’s representation from each earlier patch individually. These losses are summed across the pairs before backpropagation.
@@ 3.2 Training protocol
All models use ResNet-18 (He et al., 2016) as the backbone encoder. For action conditioning, we use a learnable linear projection to learn action embeddings (default action embedding is 128-d). The sequence aggregator in seq-JEPA is a lightweight transformer encoder (Vaswani et al., 2017) with three layers and four attention heads. The predictor is a 2-layer MLP with 1024 hidden units and ReLU activation. In order to control for and eliminate any performance gain resulting from using a transformer encoder in seq-JEPA instead of an MLP projection head, we train baselines that typically use MLP projectors in two variants: (1) with original MLP projector; and (2) with the MLP replaced by a transformer encoder and a sequence length of one. We report the better-performing variant to ensure a fair comparison. All models are trained from scratch with a batch size of 512. We use 1000 epochs for 3DIEBench and 2000 epochs for other datasets to obtain asymptotic performance. We use AdamW for models with transformer projectors (including seq-JEPA) due to its stability and improved regularization in transformer training. For ConvNet-only models with MLP heads, we use the Adam optimizer. Full hyperparameters are detailed in Appendix A.
@@ 3.3 Evaluation metrics and protocol
To assess equivariance, we follow the protocol of Garrido et al. (2023) and train a regressor on frozen encoder representations to predict the relative transformation (action) between two views. We report the $R^{2}$ score on the test set. As a proxy measure of invariance, we use top-1 classification accuracy of a linear probe on top of frozen representations. For all baselines, probes are trained on encoder outputs. For seq-JEPA, we measure accuracies on top of the aggregate representation ($z_{AGG}$ in Figure 1) and report the number of observation views used during training and inference. For completeness, we also report classification performance on top of encoder representation ($z_{i}$ in Figure 1) for seq-JEPA models in Appendix B.4. Training details of evaluation heads are given in Appendix A.2.
@@ 4.1 Quantitative evaluation on 3DIEBench
We use the 3DIEBench benchmark to quantitatively compare invariant and equivariant representations in seq-JEPA with baseline methods. This benchmark allows us to measure equivariance through decoding 3D object rotations while enabling invariance measurement through object classification. Table 3 provides a summary of our evaluation on the 3DIEBench where equivariant methods have been conditioned on rotation. In addition to the relative rotation between two views, in the last column we provide the $R^{2}$ score for predicting individual transformation parameters from representations of a single view. For seq-JEPA, we trained models with varying training sequence lengths (denoted by $M_{tr}$ in the table) and measure the linear classification performance on top of aggregate representations with different inference lengths (denoted by $M_{val}$).
Among invariant methods, BYOL achieves the highest classification accuracy, yet does not offer a high level of equivariance. Among equivariant baselines, SIE and ContextSSL yield strong rotation prediction performance due to their specialized equivariance predictors and loss functions, but underperform in classification. EquiMod and SEN offer better classification performance, yet compromise equivariance. In contrast, seq-JEPA achieves strong performance in both, matching the best rotation $R^{2}$ scores while exceeding baselines in classification, with gains increasing with inference sequence length $M_{val}$. Ablating action conditioning leads to a sharp drop in equivariance but retains classification accuracy, confirming our hypothesis that action-conditioned sequential aggregation enables a segregated invariant-equivariant representation structure. For additional results on 3DIEBench including models with varying training and inference sequence lengths and models conditioned on both rotation and color, see Appendices B.1 to B.3.
Figure 3: Top-1 linear classification (invariance) vs. rotation prediction (equivariance) performance on 3DIEBench. seq-JEPA learns good representations for both tasks. Subscripts indicate training and inference sequence lengths.
Table 1: Evaluation on 3DIEBench for linear probe classification (invariance) and rotation prediction (equivariance). Equivariant models and seq-JEPA are conditioned on rotation. For seq-JEPA, training and inference sequence lengths are denoted by $M_{tr}$ and $M_{val}$.
Method Top-1 Acc. (%) Rel. Rot. ($R^{2}$) Indiv. Rot. ($R^{2}$)
@@
seq-JEPA (1, 1) 84.08 0.65 0.69
seq-JEPA (1, 3) 85.31 0.65 0.69
seq-JEPA (3, 3) 86.14 0.71 0.74
seq-JEPA (3, 5) 87.41 0.71 0.74
seq-JEPA (no act cond) 86.05 0.29 0.37
@@ 4.2 Qualitative evaluation on 3DIEBench
To visualize equivariance in representational space, we retrieve the three nearest representations of a query image from the validation set of 3DIEBench (Figure 5). While all models retrieve the correct object category, only seq-JEPA and SIE consistently preserve rotation across all retrieved views, consistent with their high $R^{2}$ scores. Next, we projected encoder and aggregate representations using 2D UMAP (Figure 5). The left panel shows encoder representations colored by class label, while the middle panel displays the same encoder representations colored by rotation angle. The smooth color gradation across the map within each class cluster in the middle panel suggests that the encoder captures rotation angle as a continuous factor, implying equivariance to rotation (e.g., the red class in the bottom-right corner of the right panel and the corresponding part in the middle panel). The right panel shows aggregate representations colored by class label. Comparing the class-colored plots (left and right panels), we observe that both encoder and aggregate representations contain class information. However, when we aggregate multiple views of a sample, some of the intra-class variability (resulting from transformations such as rotation) is eliminated, causing each class’ representational cluster to become more homogeneous. This aggregation procedure likely reduces variation due to rotation and makes the representations more invariant, resulting in decreased intra-class spread and increased inter-class distance. We create a similar UMAP visualization for seq-JEPA with ablated rotation conditioning in Appendix B.5 to highlight the role of action conditioning in achieving equivariance to rotation.
Figure 4: Retrieval of nearest representations; given a query image, we extract the three nearest encoder representations in the validation set of 3DIEBench. The retrieved views of models with the highest quantitative rotation equivariance performance maintain the rotation of the query image across all retrieved views.
@@
Figure 5: 2-D UMAP projections of seq-JEPA’s encoder and aggregate representations trained on 3DIEBench and conditioned on rotation with $M_{tr}=3$ and $M_{val}=5$. Encoder representations for each view observation, color-coded by class (left) and rotation angle (middle). Aggregate representation for $M_{val}=5$, color-coded by class (right).
4.3 Evaluation with Hand-Crafted Augmentations
@@
We assess invariance and equivariance under hand-crafted augmentations by training on CIFAR100 and Tiny ImageNet (Table 2). Models with action conditioning are trained conditioned on crop, color jitter, blur, or all three (indicated in the first column of the table). seq-JEPA consistently achieves higher equivariance than both invariant and equivariant baselines across all transformations. Notably, except for the model trained on CIFAR-100 and conditioned on blur, the best equivariance performance for a given augmentation is achieved when the model is specialized and conditioned only on that augmentation. Furthermore, ablating actions (last row in the table) causes seq-JEPA to lose its equivariance across transformations compared to action-conditioned models. Overall, our model outperforms both invariant and equivariant families in terms of equivariance, while being competitive in terms of classification performance. For additional results with varying training and inference sequence lengths, see Appendix B.3.
@@
Table 2: Evaluation with hand-crafted augmentations on CIFAR100 and Tiny ImageNet; equivariance is measured by predicting relative transformation parameters associated with crop, color jitter, or blur augmentations. For all seq-JEPA models, $M_{val}=5$.
@@
4.4 Predictive Learning across Saccades and Path Integration
In our third action-observation setting, we consider predictive learning across simulated eye movements to exhibit seq-JEPA’s ability in leveraging a sequence of partial observations to learn visual representations in a predictive manner. Table 6 shows that seq-JEPA reaches 83.44% top-1 accuracy on STL-10, comparable to SimCLR (85.23%) trained with full-resolution images and strong augmentations. This gap narrows further when increasing the inference sequence length from $M_{val}=4$ to $M_{val}=6$. Ablating action conditioning causes the accuracy on top of the aggregate representations to drop sharply, indicating that 2-D positional awareness is essential to forming semantic representations across simulated eye movements. Compared to Conv-JEPA, which accumulates prediction losses pairwise, seq-JEPA performs better in classification, highlighting the importance of sequence aggregation when dealing with partial observations in SSL. Further ablations show that saliency-driven sampling and IoR are critical for forming informative, non-overlapping patch sequences and subsequently a high-quality aggregate representation. Interestingly, while random uniform patch sampling negatively impacts classification accuracy due to lower semantic content, it results in the highest positional equivariance as the model samples patches and corresponding saccade actions from a more diverse set of positions across the entire image, not just the salient regions. UMAP projections for PLS with and without action conditioning in Appendix B.5 further underscore the role of action conditioning in enabling positional equivariance.
Path Integration. In the context of eye movement-driven observations and in general sequential observations, another ability that naturally arises is path integration (McNaughton et al., 2006)—predicting the cumulative transformation/action from a sequence of actions. We evaluate this task in both eye movements in PLS (visual path integration) and object rotations in 3DIEBench (angular path integration). As shown in Figure 6, seq-JEPA demonstrates strong performance in both settings, with performance degrading gracefully as sequence length increases. Ablating action conditioning causes path integration to fail, whereas ablating the visual stream has only a minor impact—highlighting that action information is the dominant signal for this task. Full details of the path integration setup are provided in Appendix B.6.
Table 3: Evaluation of predictive learning across saccades on STL-10. Equivariance is measured with respect to fixation coordinates. Unless stated otherwise, $M_{tr}=M_{val}=4$.
Conditioning Top-1 Acc. (%) Position ($R^{2}$)
@@ 4.4 Predictive Learning across Saccades and Path Integration
position seq-JEPA (w/o saliency & IoR) 79.85 0.88
position seq-JEPA (w/o IoR) 77.97 0.85
@@
4.5 Ablations: Action Conditioning Is Key
To better understand the mechanisms underlying seq-JEPA’s invariant-equivariant representation learning and role of action conditioning, we perform a set of ablation experiments on 3DIEBench. Specifically, we study: (i) the role of action conditioning in the transformer and predictor; (ii) the impact of learning action embeddings and their dimensionality. Table 7 summarizes our ablation results. Removing action conditioning entirely causes a significant drop in equivariance ($R^{2}$ from 0.71 to 0.29), although classification accuracy remains high thanks to sequence aggregation and segregated invariance-equivariance in our model. Conditioning only the transformer or only the predictor leads to intermediate results, with predictor conditioning proving more critical for equivariance. We also vary the dimensionality of the learnable action embeddings: performance saturates around the default size of 128, with smaller sizes (e.g., 16 or 64) already sufficient to capture the rotation structure. Non-learnable embeddings significantly reduce equivariance, underscoring the importance of learned representations in the action pathway.
Table 4: Ablation results for action conditioning (3DIEBench). All models use $M_{tr}=3$, $M_{val}=5$ (results over three random seeds).
Variant Top-1 Acc. (%) Rotation ($R^{2}$)